Live GMAT classes with Target Test Prep start Nov 21 and Nov 22. First class free.

See classes
Target Test Prep · GMAT

Choose how you want to prepare

Learn live with an expert or move at your own pace. Every option includes the complete TTP study system.

Newsweek Readers’ Choice Awards 2026 winner: Target Test Prep, #1 Test Prep

Target Test Prep was voted #1 Best Test Prep in the 2026 Newsweek Readers’ Choice™ Awards.

★★★★★5.0563 reviews
GMATLiveTeachStarts Nov 22
Josh Braslow, Target Test Prep GMAT expert
LIVE ONLINE CLASSES

Get Ready for GMAT Test Day Faster with Live Online Classes

with Josh Braslow, 20+ years in test prep

Nov 22 · Josh Braslow
Sun · 9:30 AM to 12:30 PM ET
Included
40 hours of live online classes + 6 months of TTP OnDemand
  • Attend the first class for free
  • Every class is recorded, so you never fall behind
View classes & enroll
Limited seats availableTarget Test Prep
EALiveTeachOnDemandStart anytime
EXECUTIVE ASSESSMENT

Target Test Prep EA OnDemand

Self-paced EA prep. Study on your schedule.

EXECUTIVE ASSESSMENT

with

165+ EA score guarantee
$05-day trial no automatic billing
Schedule
Included
40 hours of live online classes plus six months of access to the complete TTP EA OnDemand course.
  • 165+ EA Score Guarantee
  • 4,100+ Quant, Verbal, and Integrated Reasoning practice questions
  • 400+ hours of in-depth video lessons
  • 3,000+ step-by-step video solutions
View EA class & enroll Start free 5-day trial
Limited cohort · enrollment openTrial includes full course accessTarget Test Prep
GMATOnDemandStart anytime
SELF-PACED MASTERCLASS

Target Test Prep GMAT OnDemand

Complete access from day one. Study on your schedule.

715+ score guarantee
$0to start then $127/mo
  • Personalized study plan and analytics
  • Thousands of lessons and practice questions

Probability Problem

Expert replies
by yuliawati » Wed Oct 27, 2010 10:26 pm
How to solve this problem quickly? I need more than 5 mins to get the answer. A long calculation.... :-(
Source: MGMAT

Bill and Jane play a simple game involving two fair dice, each of which has six sides numbered from 1 to 6 (with an equal chance of landing on any side). Bill rolls the dice and his score is the total of the two dice. Jane then rolls the dice and her score is the total of her two dice. If Jane’s score is higher than Bill’s, she wins the game. What is the probability the Jane will win the game?

(A) 15/36
(B) 175/432
(C) 575/1296
(D) 583/1296
(E) 1/2

OA later
Join the discussion
Source: — Problem Solving |

by Rezinka » Wed Oct 27, 2010 11:09 pm
Is OA E?
Join the discussion

by yuliawati » Thu Oct 28, 2010 2:42 am
No, the OA isn't E.
Join the discussion

by Rahul@gurome » Thu Oct 28, 2010 3:27 am
yuliawati wrote:How to solve this problem quickly? I need more than 5 mins to get the answer. A long calculation.... :-(
Source: MGMAT

Bill and Jane play a simple game involving two fair dice, each of which has six sides numbered from 1 to 6 (with an equal chance of landing on any side). Bill rolls the dice and his score is the total of the two dice. Jane then rolls the dice and her score is the total of her two dice. If Jane’s score is higher than Bill’s, she wins the game. What is the probability the Jane will win the game?

(A) 15/36
(B) 175/432
(C) 575/1296
(D) 583/1296
(E) 1/2

OA later
Yes, this problem involves a long calculation to some extent.
But we can reduce the calculations to some extent by noting that Bill and Jane has equal probability to win the game. But that doesn't mean that the probability is 1/2, as there also exists the case of tie.

Thus, the required probability = [1 - Probability of tie]/2

Now the possible scores are,
2 : In 1 way => (1, 1)
3 : In 2 way => (1, 2), (2, 1)
4 : In 3 way => (1, 3), (2, 2), (3, 1)
5 : In 4 way => (1, 4), (2, 3), (3, 2), (4, 1)
6 : In 5 way => (1, 5), (2, 4), (3, 3), (4, 2), (5, 1)
7 : In 6 way => (1, 6), (2, 5), (3, 4), (4, 3), (5, 2), (6, 1)
8 : In 5 way => (2, 6), (3, 5), (4, 4), (5, 3), (6, 2)
9 : In 4 way => (3, 6), (4, 5), (5, 4), (6, 3)
10 : In 3 way => (4, 6), (5, 5), (6, 4)
11 : In 2 way => (5, 6), (6, 5)
12 : In 1 way => (6, 6)

Now, Probability of tie
= Probability that scores of Bill and Jane is same
= P(2) + P(3) + P(4) + P(5) + P(6) + P(7) + P(8) + P(9) + P(10) + P(11) + P(12)
= (1 + 4 + 9 + 16 + 25 + 36 + 25 + 16 + 9 + 4 + 1)/(36*36)
= 146/1296

Thus, required probability = [1 - (146/1296)]/2 = 575/1296

The correct answer is C.
Rahul Lakhani
Quant Expert
Gurome, Inc.
https://www.GuroMe.com
On MBA sabbatical (at ISB) for 2011-12 - will stay active as time permits
1-800-566-4043 (USA)
+91-99201 32411 (India)
Join the discussion

by GMATGuruNY » Thu Oct 28, 2010 3:36 am
yuliawati wrote:How to solve this problem quickly? I need more than 5 mins to get the answer. A long calculation.... :-(
Source: MGMAT

Bill and Jane play a simple game involving two fair dice, each of which has six sides numbered from 1 to 6 (with an equal chance of landing on any side). Bill rolls the dice and his score is the total of the two dice. Jane then rolls the dice and her score is the total of her two dice. If Jane’s score is higher than Bill’s, she wins the game. What is the probability the Jane will win the game?

(A) 15/36
(B) 175/432
(C) 575/1296
(D) 583/1296
(E) 1/2

OA later
3 possible outcomes:
P(J) = Probability that Jane wins
P(B) = Probability that Bill wins
P(D) = Probability in a draw (each gets the same score)

P(J) + P(B) + P(D) = 1.
Since Jane and Bill each have the same chance of winning, P(J) = P(B).
Thus:
P(J) + P(J) + P(D) = 1
2*P(J) + P(D) = 1
P(J) = [1- P(D)] / 2

So to determine P(J), we need to determine P(D).

The probabilities of getting the different sums display the following symmetry:

P(2) = P(12)
P(3) = P(11)
P(4) = P(10)
P(5) = P(9)
P(6) = P(8)

Let's determine the various probabilities to discern the pattern:
P(2) = P(1,1) = 1/6 * 1/6 = 1/36.
P(3) = P(1,2) + P(2,1) = 1/36 + 1/36 = 2/36
P(4) = P(1,3) + P(3,1) + P(2,2) = 3/36
Now we see the pattern:
P(5) = 4/36
P(6) = 5/36
P(7) = 6/36

A draw means that each gets the same score:
P(both score 2) = P(2) * P(2) = 1/36 * 1/36 = 1/1296
P(both score 3) = P(3) * P(3) = 2/36 * 2/36 = 4/1296
P(both score 4) = 3/6 * 3/36 = 9/1296
P(both score 5) = 4/36 * 4/36 = 16/1296
P(both score 6) = 5/36 * 5/36 = 25/36

P(both score 2 through 6) = (1+4+9+16+25)/1296 = 55/1296
Thus, P(both score 8 through 12) = 55/1295
P(both score 7) = 6/36 * 6/36 = 36/1296
Thus, P(D) = (55+55+36)/1296 = 146/1296

Thus P(J) = (1 - 146/1296)/2 = 575/1296.

The correct answer is C.
Last edited by GMATGuruNY on Sun Feb 26, 2012 5:08 am, edited 2 times in total.
Private tutor exclusively for the GMAT and GRE, with over 20 years of experience.
Followed here and elsewhere by over 1900 test-takers.
I have worked with students based in the US, Australia, Taiwan, China, Tajikistan, Kuwait, Saudi Arabia -- a long list of countries.
My students have been admitted to HBS, CBS, Tuck, Yale, Stern, Fuqua -- a long list of top programs.

As a tutor, I don't simply teach you how I would approach problems.
I unlock the best way for YOU to solve problems.

For more information, please email me (Mitch Hunt) at [email protected].
Student Review #1
Student Review #2
Student Review #3
Join the discussion

by Rezinka » Thu Oct 28, 2010 3:37 am
Great..
Thanks
Join the discussion

by yuliawati » Thu Oct 28, 2010 3:49 am
Thank you, Rahul and Mitch!
In dealing with the probability problem, my big mistake is careless calculation. That's very easy to make a mistake^^
Join the discussion